# Non-ultralocal classical $r$-matrix structure

### v2

@@ 1 Introduction
with spectral parameter $z$ [22]. Then ${\rm tr}L^{k}(z)$ are generating functions (in $z$) of conservation laws, and they are in involution with respect to the Poisson brackets (1.2) $\{{\rm tr}L^{k}(z),{\rm tr}L^{m}(w)\}=0$ due to existence of the classical $r$-matrix [35, 13]:
@@ 2.2 Spin and multispin Calogero-Moser model
$\begin{array}[]{c}\displaystyle{\{U_{1}(z,x),U_{2}(w,y)\}=-k\sum\limits_{i=1}^{N}E_{ii}\otimes E_{ii}\Big{(}E_{1}(z)+E_{1}(w)\Big{)}\delta^{\prime}(x-y)+}\\ \displaystyle{+\sum\limits_{i\neq j}^{N}E_{ii}\otimes E_{ij}\,f(w,q_{ij}(x))\alpha_{j}^{2}(x)\delta(x-y)-\sum\limits_{i\neq j}^{N}E_{ij}\otimes E_{ii}\,f(z,q_{ij}(x))\alpha_{j}^{2}(x)\delta(x-y)+}\\ \displaystyle{+\sum\limits_{i\neq j}^{N}E_{ij}\otimes E_{jj}\Big{(}f(z,q_{ij}(x))\alpha_{j}^{2}(x)\delta(x-y)-\phi(z,q_{ij}(x))k\delta^{\prime}(x-y)\Big{)}-}\\ \displaystyle{-\sum\limits_{i\neq j}^{N}E_{ii}\otimes E_{ji}\Big{(}f(z,q_{ji}(x))\alpha_{i}^{2}(x)\delta(x-y)+\phi(z,q_{ji}(y))k\delta^{\prime}(x-y)\Big{)}\,.}\end{array}$ (3.22)
@@
$\begin{array}[]{c}\displaystyle{f(z,q_{ji}(x))\alpha_{i}^{2}(x)\delta(x-y)+\phi(z,q_{ji}(y))k\delta^{\prime}(x-y)=}\\ \\ \displaystyle{=f(z,q_{ji}(x))\alpha_{i}^{2}(x)\delta(x-y)+\phi(z,q_{ji}(x))k\delta^{\prime}(x-y)+k(q_{j,x}-q_{i,x})f(z,q_{ji}(x))\delta(x-y)=}\\ \\ \displaystyle{=f(z,q_{ji}(x))\alpha_{j}^{2}(x)\delta(x-y)+\phi(z,q_{ji}(x))k\delta^{\prime}(x-y)\,.}\end{array}$ (3.23)
@@
$\begin{array}[]{c}\displaystyle{\mu_{i}=-c\,,\quad i=1,...,N\,,}\end{array}$ (3.40)
@@ 6 Appendix B: elliptic $R$-matrices
It is deduced from the quantum Yang-Baxter equation (B.2) in the limit (B.8).

### v3

@@ 1 Introduction
with spectral parameter $z$ [22]. Then ${\rm tr}L^{k}(z)$ are generating functions (in $z$) of conservation laws, and they are in involution with respect to the Poisson brackets (1.2) $\{{\rm tr}L^{k}(z),{\rm tr}L^{m}(w)\}=0$ due to existence of the classical $r$-matrix [35, 13] [^1] Let us mention that for skew-symmetric $r$-matrices (1.4) takes the form $\{L_{1}(z),L_{2}(w)\}=[L_{1}(z)+L_{2}(w),r_{12}(z,w)]$. The classical $r$-matrices under consideration, interrelations between them and the corresponding Yang-Baxter equations are reviewed in the Appendix B.:
@@ 2.2 Spin and multispin Calogero-Moser model
$\begin{array}[]{c}\displaystyle{\{U_{1}(z,x),U_{2}(w,y)\}=-k\sum\limits_{i=1}^{N}E_{ii}\otimes E_{ii}\Big{(}E_{1}(z)+E_{1}(w)\Big{)}\delta^{\prime}(x-y)+}\\ \displaystyle{+\sum\limits_{i\neq j}^{N}E_{ii}\otimes E_{ij}\,f(w,q_{ij}(x))\alpha_{j}^{2}(x)\delta(x-y)-\sum\limits_{i\neq j}^{N}E_{ij}\otimes E_{ii}\,f(z,q_{ij}(x))\alpha_{j}^{2}(x)\delta(x-y)+}\\ \displaystyle{+\sum\limits_{i\neq j}^{N}E_{ij}\otimes E_{jj}\Big{(}f(z,q_{ij}(x))\alpha_{j}^{2}(x)\delta(x-y)-\phi(z,q_{ij}(x))k\delta^{\prime}(x-y)\Big{)}-}\\ \displaystyle{-\sum\limits_{i\neq j}^{N}E_{ii}\otimes E_{ji}\Big{(}f(w,q_{ji}(x))\alpha_{i}^{2}(x)\delta(x-y)+\phi(w,q_{ji}(y))k\delta^{\prime}(x-y)\Big{)}\,.}\end{array}$ (3.22)
@@
$\begin{array}[]{c}\displaystyle{f(w,q_{ji}(x))\alpha_{i}^{2}(x)\delta(x-y)+\phi(w,q_{ji}(y))k\delta^{\prime}(x-y)=}\\ \\ \displaystyle{=f(w,q_{ji}(x))\alpha_{i}^{2}(x)\delta(x-y)+\phi(w,q_{ji}(x))k\delta^{\prime}(x-y)+k(q_{j,x}-q_{i,x})f(w,q_{ji}(x))\delta(x-y)=}\\ \\ \displaystyle{=f(w,q_{ji}(x))\alpha_{j}^{2}(x)\delta(x-y)+\phi(w,q_{ji}(x))k\delta^{\prime}(x-y)\,.}\end{array}$ (3.23)
@@
$\begin{array}[]{c}\displaystyle{\mu_{i}=c\,,\quad i=1,...,N\,,}\end{array}$ (3.40)
@@ 6 Appendix B: elliptic $R$-matrices
It is deduced from the quantum Yang-Baxter equation (B.2) in the limit (B.8). Let us mention that from the classical mechanics viewpoint the Yang-Baxter equation (B.9) is a sufficient condition for the Jacobi identity $\{\{L_{1}(z_{1}),L_{2}(z_{2})\},L_{3}(z_{3})\}+{\rm cycl.}=0$ for non-dynamical skew-symmetric (i.e. $r_{12}(z_{1},z_{2})=-r_{21}(z_{2},z_{1})$) $r$-matrix and the classical $r$-matrix structure
$\begin{array}[]{c}\displaystyle{\{L_{1}(z_{1}),L_{2}(z_{2})\}=[L_{1}(z_{1})+L_{2}(z_{2}),r_{12}(z_{1},z_{2})]\,.}\end{array}$ (B.10)
In the general case (when non-dynamical $r$-matrix is not skew-symmetric) the classical exchange relation is
$\begin{array}[]{c}\displaystyle{\{L_{1}(z_{1}),L_{2}(z_{2})\}=[L_{1}(z_{1}),r_{12}(z_{1},z_{2})]-[L_{2}(z_{2}),r_{21}(z_{2},z_{1})]}\end{array}$ (B.11)
and the classical Yang-Baxter equation has the form:
$\begin{array}[]{c}\displaystyle{[r_{12}(z_{1},z_{2}),r_{13}(z_{1},z_{3})]+[r_{12}(z_{1},z_{2}),r_{23}(z_{2},z_{3})]+[r_{32}(z_{3},z_{2}),r_{13}(z_{1},z_{3})]=0\,.}\end{array}$ (B.12)
For skew-symmetric non-dynamical $r$-matrices (B.12) turns into (B.9) and (B.11) turns into (B.10).